The High Altitude Ski Company manufactures skis in the finishing and assembly process. Skis are manufactured in 50-ski batch sizes. The finishing time is 2 minutes per ski. The assembly time is 7 minutes per ski. It takes 18 minutes to move a batch of skis from finishing to assembly.

a. Compute the value-added, non-value-added, and total lead time of this process.
Answer #1

Value added lead time = Finishing time + Assembly time = 2 + 7 = 9 minutes

Non-value added lead time = [Value added time x (Batch size - 1)] + Move-in time = [9 x (50 - 1) + 18 = 459 minutes

Total lead time = Value added lead time + Non-value added lead time = 9 + 459 = 468 minutes